OpenStudy (anonymous):

What is the determinator of (1/2-2/3)+(1/2+1/3)

OpenStudy (anonymous):

do u have any idea?

OpenStudy (anonymous):

No,and Its a fraction

OpenStudy (anonymous):

Do you mean the denominator?

OpenStudy (anonymous):

Yeah denominator..my bad..

OpenStudy (anonymous):

It's cool

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

And actually I'm sorry because I don't really know this stuff :/

OpenStudy (anonymous):

if 1/2 * 3 = 3/6

OpenStudy (anonymous):

well now I know the answer

OpenStudy (anonymous):

-2/3 *2 = -4/6

OpenStudy (anonymous):

now u can subtract the number 3/6-4/6 = ??

OpenStudy (anonymous):

The answe should be a whole number

OpenStudy (anonymous):

no

OpenStudy (anonymous):

You calculate the problem then reduced answer that u get to simplest terms. so what is the denominator of the resulting fraction